I 22. Which gas would effuse twice as fast as sulfur dioxide (SO2)? Не А. CH4 В. O2 С. PF5 D. XeF6 Е.
Answer #1

Rate of effusion is inversely proportional to square root of molar mass

rate = k*sqrt(1/M)

M is molar mass

If we take ratio, above expression becomes

rate(SO2)/rate(unknown) = sqrt (M(unknown)/M(SO2))

we have:

M(SO2) = 64.07 g/mol

rate(SO2) = 1.0

rate(unknown) = 2.0

rate(SO2)/rate(unknown) = sqrt (M(unknown)/M(SO2))

1.0/2.0 = sqrt (M(unknown)/64.07)

sqrt (M(unknown)/64.07) = 0.5

(M(unknown)/64.07) = 0.25

M(unknown) = 16.02 g/mol

Molar mass is 16.02 g/mol for CH4

Answer: B